Which of the numbers 32,23,54,and 37 are prime?
Answer by jim_thompson5910(35256)

Solved by pluggable solver: Find factors of any number
32 is NOT a prime number: 32 = 2 * 2 * 2 * 2 * 2

Work Shown

32 is divisible by 2: 32 = 16 * 2.
16 is divisible by 2: 16 = 8 * 2.
8 is divisible by 2: 8 = 4 * 2.
4 is divisible by 2: 4 = 2 * 2.
2 is not divisible by anything.

This calculation used this Prime Factorization Algorithm.



Solved by pluggable solver: Find factors of any number
23 is a prime number

Work Shown

23 is not divisible by anything.

This calculation used this Prime Factorization Algorithm.



Solved by pluggable solver: Find factors of any number
54 is NOT a prime number: 54 = 2 * 3 * 3 * 3

Work Shown

54 is divisible by 2: 54 = 27 * 2.
27 is divisible by 3: 27 = 9 * 3.
9 is divisible by 3: 9 = 3 * 3.
3 is not divisible by anything.

This calculation used this Prime Factorization Algorithm.



Solved by pluggable solver: Find factors of any number
37 is a prime number

Work Shown

37 is not divisible by anything.

This calculation used this Prime Factorization Algorithm.




So the only prime numbers are 23 and 37